OpenStudy (anonymous):

Find the equation of the circle with center at (3, 2) and through the point (5, 4).

OpenStudy (usukidoll):

ponies!!! :D

OpenStudy (anonymous):

hey! hot Rarity :D lol but i really need an answer

OpenStudy (usukidoll):

I only know part of it :(

OpenStudy (usukidoll):

I haven't touch this problem in eons x.x

OpenStudy (usukidoll):

x^2+y^2 = 1 is the default equation of the circle. the center is (0,0) and the radius is one.

OpenStudy (usukidoll):

since the center is at 3,2 it's (x-3)^2+(y-2)^2 but to find the radius.. x.x.

OpenStudy (usukidoll):

D: arghhh don't use the math you lose the math sad bambi@!

OpenStudy (anonymous):

@Kainui

OpenStudy (usukidoll):

D: I can't help it if higher math courses only use some topics DKKKK

OpenStudy (usukidoll):

Elementary Linear Algebra had nothing to do with this D:

OpenStudy (usukidoll):

this shows up again in Calculus III, but I had the simple questions x.x

OpenStudy (fifciol):

find radius \[r^2= (x_b-x_a)^2+ (y_b-y_a)^2\]

OpenStudy (zzr0ck3r):

to find the radius \[\sqrt{(5-3)^2(4-2)^2}=r\]

OpenStudy (zzr0ck3r):

so r = err add a plus

OpenStudy (zzr0ck3r):

\[\sqrt{(5-3)^2+(4-2)^2}=r\]

OpenStudy (anonymous):

i got 6

OpenStudy (zzr0ck3r):

\[r=2\sqrt{2}\\so\\(x-3)^2+(2-y)^2=2\sqrt{2}\]

OpenStudy (fifciol):

nope, in equation of a circle is r^2

OpenStudy (fifciol):

so it schould be 8

OpenStudy (fifciol):

instead of 2sqrt2

OpenStudy (zzr0ck3r):

\[(x-3)^2+(y-2)^2=(2\sqrt{2})^2=8\]

OpenStudy (zzr0ck3r):

sorry its late

OpenStudy (usukidoll):

DISTANCE FORMULA UGHHHHHHHHHH! *bangs head* of course! Use that and then square it... gawddd D((((

OpenStudy (anonymous):

so then the equation of a circle that passes through those points is...?

OpenStudy (zzr0ck3r):

\[(x-3)^2+(y-2)^2=8\]

OpenStudy (usukidoll):

yuppppp